# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ 6       Solve the following optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}}\left\{\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda N_{k}}\right\}$ (3)
@@
Algorithm 2 Adaptive Policy Learning via Adjusted Logarithmic Smoothing
Note that at time $k$, we instantiate $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})$ with $(\pi_{j})_{j=0\cdots k}$ the previous outputs of Algorithm 2.
Note that in Thm. 4.1, contrary to Sec. 3.2, we do not control directly $R(\pi_{Q})$ but the difference $R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$. This means that running Algorithm 2 at time $k$ using $\cup_{j=0}^{k}S_{j}$ tends to ensure a better generalisation ability than what we obtained (on a weighted average) by running Algorithm 2 on $\cup_{j=0}^{k-1}S_{j}$. This phenomenon gives an intuition about why Algorithm 2 converge quickly to the optimum. We prove this intuition in what follows.
Towards an accelerated convergence rate for Algorithm 2. Recall that $\Pi$ is the considered space of policies on which the optimization of the risk is conducted. If $\Pi$ is the space of all possible policies, the optimal policy $\pi_{\star}$ can be characterized. For any context $x$, let $a_{\star}(x)=\operatorname*{arg\,min}_{a^{\prime}\in\mathcal{A}}c(a^{\prime},x)$ with ties broken arbitrarily. We obtain the optimal policy $\pi_{\star}$ as the one that chooses deterministically the optimal action $a_{\star}(x)$ for any context $x$, giving:
In the following, we assume that our class of parametrized predictors $\{f_{\theta}\mid\theta\in\Theta\}$ is rich enough to represent $\pi_{\star}$. Then, there exists a $Q^{\star}\in\mathcal{C}(\Theta)$ such that $\pi_{\star}=\pi_{Q^{\star}}$. To prove the fast convergence of Algorithm 2 to $\pi_{\star}$, we state a first proposition.
@@
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$0\leq R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{1-\lambda}\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}},$
with $L(\pi,\pi_{j})=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\left(\pi(a|x)\left(\left(\frac{1}{\pi_{j}(a|x)-1}\right)^{2}-1\right)+1\right)c^{2}\right]\,.$
Proposition 4.2 is proven in Appendix E and highlights the interest of $h^{\text{adj}}_{\lambda}$: we avoid the pseudo variance terms $\mathcal{S}_{j}(\pi_{\star})$ of Algorithm 1 to get $L\left(\pi_{\star},\pi_{j}\right)$. Contrary to $S_{j}(\pi_{\star})$, we have that, for deterministic policies $\pi_{\star}$, $L(\pi_{\star},\pi_{j})\rightarrow 0$ if $\pi_{j}\rightarrow\pi_{\star}$. Such desirable behavior leaves room for accelerated convergence rate. To investigate this route, we target the convergence towards the optimal policy $\pi_{\star}$ over the whole space and study the behavior of $L$.
To prove the acceleration of our method, we need to define additional quantities. We define the action $a$ sub-optimality given a context $x$ by: $\Delta_{a,x}:=c(a,x)-c(a_{\star}(x),x)\geq 0$. Then, for any $u\in[0,1)$, we define $\Delta_{u}$ such that:
@@
For a fixed $u$, $\Delta_{u}$ quantifies how separable $a^{\star}(x)$ is from the rest of actions. In other words, the magnitude of $\Delta_{u}$ quantifies the difficulty of the problem. $\Delta_{u}=0$ for all $u\in[0,1)$ means that the optimal action $a^{\star}(x)$ is not distinguishable from the second best action, almost surely. Oppositely, $\Delta_{u}=1$ for a specific $u$ means that with probability $1-u$, the costs are deterministic and the optimal action $a^{\star}(x)$ is the only one to have a non null cost (of $-1$). We want to avoid the first case, as the existence of a strictly positive $\Delta_{u}>0$ will be important next. Finally, we introduce the minimal plausibility of the optimal actions under $\pi_{k}$:
@@
Proof is in Appendix E. The existence of a $\Delta_{u}>0$ is similar to margin conditions used in multi-arm contextual bandits fast_rate_cb ; adaptive , and $C^{\star}_{k}>0$ is a condition on the coverage of the optimal actions by $\pi_{k}$. In our adaptive algorithm for instance, $C^{\star}_{k}$ will intuitively grow as $\pi_{k}$ gets closer to $\pi_{\star}$. Note that the condition $C^{\star}_{k}$ is less restrictive than the uniform coverage assumption used in adaptive ; scrm . A similar inequality on the variance of importance weights was used as an assumption to obtain fast rates in scrm . Here, we prove such result and characterize when it is possible to obtain. For instance, this inequality cannot hold when the costs are constant. Finally, we used $\Delta_{u}$ and $C^{\star}_{k}$ for the sake of interpretability, and give a sharper, yet less interpretable condition in Sec. E.3.
Lemma 4.3 shows that the expansion of $L$ is controlled by a difference of risks. Smaller $\gamma_{k}$ sharpens this inequality, which is the case for problems with distinguishable optimal actions and near optimal logging policies $\pi_{k}$. Plugging this inequality in Proposition 4.2 then yields a recursive control on the sequence $\left(R(\pi_{k})\right)_{k\geq 0}$ (see Cor. E.5). Such a recursive relation can then be transformed in to an explicit accelerated convergence rate for the output of Algorithm 2 at the cost of the following mild assumption, where we suppose that our learning algorithm outputs for all $k\geq 0$, $\pi_{k}$ such that:
$\forall k\geq 0\,,\quad\gamma_{k}\leq\gamma<+\infty.$
This holds when $C^{\star}_{k}\geq C^{\star}>0$, which is a reasonable assumption as we expect $\pi_{0}$ to cover optimal actions and $\pi_{k},k\geq 1$ to be closer to $\pi_{\star}$. With this, we state our principal result below.
@@
Let $n_{j}=m>0$ for all $j$. Then, for any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm 2 at time $k$ with parameter $\lambda_{m}:=\frac{1-\alpha}{8\gamma\sqrt{m}}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
@@
Thm. 4.4 is proven in Appendix E and provides an accelerated rate of $\mathcal{O}\left(\frac{1}{\sqrt{m}k^{\alpha}}\right)$ compared to the $\mathcal{O}\left(\frac{1}{\sqrt{mk}}\right)$ of Algorithm 1 as long as $\alpha>1/2$. Again, we recover the same rate than sakhi2024logarithmic in the batch setting (i.e. policy after one batch update: $k=1,m=n$) and improve the rate in the online setting (i.e. updating the behavior policy after each interaction: $k=n,m=1$) to $\mathcal{O}(1/n^{\alpha})$. Thm. 4.4 also works for any intermediate regime between batch and online settings, and suggests that considering several updates is beneficial to learn efficiently. Indeed, as long as $\alpha>1/2$, the number of updates $k$ will have a stronger influence than the size of the batch $m$. Note however that $C_{\alpha}$, the constant behind is deteriorating as $\alpha$ is close to $1$. Bigger values of $\lambda_{m}$ result in smaller constants but slower convergence rates, while smaller $\lambda_{m}$ trade a heavy constant for a fast convergence rate. There is then a practical balance to find between a fast convergence rate and reasonable values of $C_{\alpha}$.
Finally, to tackle a large-range of real-life scenarios, where $n_{k}$ may vary through time, we provide additional convergence rates Algorithm 2 encompassing this setting in Cor. 4.5.
@@
Let $\beta_{1}:=\sup_{k\geq 0}\frac{k+1}{N_{k}}\leq 1$. Assume there exists $\beta_{2}\in\mathbb{R}^{+}$ such that for all $k$, $n_{k}\leq\beta_{2}$. For any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm 2 at time $k$ with parameter $\lambda=\frac{1}{1+2^{2+\alpha}\gamma\beta_{1}\beta_{2}B(\alpha)}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
$R(\pi_{k})-R(\pi^{\star})\leq\frac{C_{\alpha}}{(k+1)^{\alpha}},$
where $C_{\alpha}\leq\max\left(1,\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda^{2}\gamma\beta_{2}B(\alpha)}\right)$, and $B(\alpha)=\frac{1}{1-\alpha}$.
Observe that the convergence rate given in Cor. 4.5 is weaker than that of Thm. 4.4, as it does not directly depend on the batch size $n_{j}$. This relaxation is necessary to accommodate adversarial scenarios where $n_{j}$ can be as small as 1. For instance, consider $n\geq 2$ and a sequence such that for all $j\geq 1$, $n_{2j-1}=m$ and $n_{2j}=1$. In this case, for all $k\geq 0$, we have $N_{2k-1}=k(m+1)-1$ and $N_{2k}=k(m+1)$. Thus, for all $k$, it holds that $\frac{k}{N_{k}}\leq\frac{2}{m}$, and the assumptions of Cor. 4.5 are satisfied. Importantly, Cor. 4.5 matches the optimal convergence rate of Thm. 4.4 in the fully online regime (i.e., $k=n$, $n_{j}=1$). Having established that Algorithm 2 enjoys provably accelerated convergence rates, we turn to empirical validation to assess whether this theoretical advantage is reflected in practice.
@@ 5 Experiments
General Setup. Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work (dudik14doubly, ; swaminathan2015batch, ). Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies sakhi2023pac ; aouali23a , these policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost. Appendix F details the experimental design and obtained results.
Effect of Multiple Deployments. Unlike prior batch approaches, our adaptive algorithms support multiple updates of the behavior policy $\pi_{0}$, while seamlessly reducing to the batch case when $k=1$. Using the MNIST and CIFAR100 datasets, we evaluate how the number of intermediate policy updates $k$ affects the performance of the final learned policy $\pi$. Starting from an approximately uniform policy $\pi_{0}$, we run both adaLS (Alg 1) and the improved adaAdjLS (Alg 2). For each $k\in\{1,5,10,100\}$, we fix the total number of interactions to $N_{k}=N$, resulting in $m=N/k$ interactions per deployment. The case $k=1$ corresponds to the static setting studied in sakhi2024logarithmic , where all data is collected using $\pi_{0}$. For $k>1$, each intermediate policy $\pi_{j}$ interacts with $m$ examples, collectively producing $N$ adaptively collected samples. We set $\lambda\sim\mathcal{O}(1/\sqrt{m})$, following our theoretical guidance. Table 1 summarizes the results: increasing $k$ consistently improves final policy performance, with diminishing returns as $\pi$ nears optimality (e.g., MNIST at $k=10$ and $100$). The gains from adaptivity are especially notable on CIFAR100, where the larger action space ($|\mathcal{A}|=100$) makes uniform policies particularly suboptimal, thus highlighting the importance of the adaptive updates. Finally, our results demonstrate the empirical advantage of adaAdjLS, validating the need for our refined estimator. Across all settings, and especially the more challenging ones, adaAdjLS consistently outperforms adaLS, providing strong empirical evidence that the adjustment is essential for achieving accelerated convergence.
@@
Dataset $k$ adaLS (Alg 1) adaAdjLS (Alg 2)
MNIST $1$ $-0.674$ $-0.678$
$5$ $-0.758$ $-0.766$
$10$ $-0.782$ $\mathbf{-0.791}$
$100$ $-0.781$ $\mathbf{-0.789}$
CIFAR100 $1$ $-0.197$ $-0.207$
$5$ $-0.344$ $-0.380$
$10$ $-0.406$ $-0.455$
$100$ $-0.508$ $\mathbf{-0.558}$
Figure 1: $R(\pi_{k})$ at each intermediate step, for $k=10$.
Benchmarks. Having demonstrated the benefits of intermediate updates and the effectiveness of our proposed adaAdjLS algorithm, we now compare our methods against the recently introduced SCRM approach scrm , which extends the CRM principle of swaminathan2015batch to the adaptive setting. We evaluate all methods on MNIST, FashionMNIST, EMNIST and CIFAR100 of different action space sizes, fixing the total number of interactions to $N$, setting $k=10$, and using theoretically recommended values for $\lambda$ for a fair comparison. adaLS and adaAdjLS divide the interaction budget uniformly across rounds ($n_{j}=N/k$), reusing all previously collected data to update policies. In contrast, SCRM employs exponentially increasing interaction sizes ($n_{j}=n_{0}\cdot 2^{j}$ with $n_{0}=\lceil N/2^{k}\rceil$) and updates policies using only the most recent batch, rather than the cumulative data ($n_{j}$ instead of $N_{j}$). Each method is run across six random seeds, and the results are summarized in Fig. 1. We find that SCRM tends to be overly conservative, particularly in early rounds where the number of interactions is small. While it does improve upon the initial behavior policy $\pi_{0}$ over time, it lags significantly behind both of our adaptive methods. Although it narrows the gap with adaLS on CIFAR100, it remains clearly outperformed—especially by adaAdjLS, which consistently achieves superior results across all datasets. These findings further confirm the empirical advantages of our approach. Additional experimental details and results can be found in Appendix F.

### v2

@@ 6   Solve the following optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}}\left\{\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda N_{k}}\right\}$
@@
Algorithm 2 Sequential Policy Learning via Adjusted Logarithmic Smoothing
While Sec.˜3.2 provides a natural sequential extension of Sakhi et al. (2024), Algorithm˜1 still suffers (Algorithm˜1) from a slow convergence rate of $\mathcal{O}(1/\sqrt{km})$ due to the nature of the performance objective, and consequently the pseudo-variance term. Is it then possible to avoid such a term? The answer is yes, by adjusting $h^{\lambda-\mathrm{Lin}}$ introduced in Sec.˜3: we define $h^{\text{adj}}_{\lambda}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q(1+\lambda c)}\right)$ and denote it as the adjusted linear LS regularizer. We then define the adjusted risk for any sequential algorithm Alg and $\lambda\in(0,1),k\geq 0$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi)$ $:=R_{0,k}^{h^{\text{adj}}_{\lambda}}(\pi)$
$=-\frac{1}{N_{k}}\sum_{j=0}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\lambda}$
$\quad\cdot\log\left(1-\frac{\lambda c_{i}}{\pi_{j}(a_{i}|x_{i})(1+\lambda c_{i})}\right)\,$,
where again, $\pi_{k+1}$ is the output of Alg at time $k$. Using similar tools to those in Sec.˜3.2, we are able to derive the following theoretical guarantee for our novel adjusted regularizer $h^{\text{adj}}_{\lambda}$.
Theorem 4.1.
For any sequential algorithm Alg, any sequence $(n_{k})_{k\geq 0}$ and data independent prior $P$. For any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)\\ \leq\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}},$
The proof of Thm.˜4.1 is established in Appendix˜D. Thm.˜4.1 relies on the same analytical tools as Sec.˜3.2, and therefore inherits the same general insights. Notably, the result remains valid for varying batch sizes $(n_{k})_{k\geq 0}$, which makes it applicable to real-world scenarios where data may arrive in irregular quantities. This theorem serves as the theoretical foundation of the main contribution of this work: a sequential learning algorithm for off-policy contextual bandits, presented in Algorithm˜2.
Note that at time $k$, we instantiate $\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})$ with $(\pi_{j})_{j=0\cdots k}$ the previous outputs of Algorithm˜2.
Note that in Thm.˜4.1, contrary to Sec.˜3.2, we do not control directly $R(\pi_{Q})$ but the difference $R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$. This means that running Algorithm˜2 at time $k$ using $\cup_{j=0}^{k}S_{j}$ tends to ensure a better generalisation ability than what we obtained (on a weighted average) by running Algorithm˜2 on $\cup_{j=0}^{k-1}S_{j}$. This phenomenon gives an intuition about why Algorithm˜2 converges faster to the optimum. We prove this intuition in what follows.
Accelerated convergence rate for Algorithm˜2. To prove the fast convergence of Algorithm˜2 to $\pi_{\star}$, we state a first proposition.
@@
For any data independent prior $P$, any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$, we have for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Algorithm˜2 at time $k$:
$0\leq R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)\\ \leq\frac{\lambda}{1-\lambda}\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}L\left(\pi_{\star},\pi_{j}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda N_{k}},$
where for all $\pi$,
$L\left(\pi,\pi_{j}\right)=\mathbb{E}_{x\sim\nu}\left[\mathbb{E}_{\pi_{j}}\left[c^{2}\right]+\mathbb{E}_{\pi}\left[c^{2}\left(\frac{1}{\pi_{j}(a|x)}-2\right)\right]\right]$.
Proposition˜4.2 is proven in Appendix˜D and highlights the interest of $h^{\text{adj}}_{\lambda}$: we avoid the pseudo variance terms $\mathcal{S}_{j}(\pi_{\star})$ of Algorithm˜1 to get $L\left(\pi_{\star},\pi_{j}\right)$. Contrary to $S_{j}(\pi_{\star})$, we have that, for any deterministic policy $\pi_{d}$, $L(\pi_{d},\pi_{j})\rightarrow 0$ if $\pi_{j}\rightarrow\pi_{d}$. Such desirable behavior leaves room for accelerated convergence rates. To investigate this route, we target the convergence towards the optimal policy $\pi_{\star}$ and study the behavior of $L$.
To prove the acceleration of our method, we need to define additional quantities. We define the sub-optimality of action $a$, given a context $x$ by: $\Delta_{a,x}:=c(a,x)-c(a_{\star}(x),x)\geq 0$. Then, for any $u\in[0,1)$, we define $\Delta_{u}$ such that:
@@
For a fixed $u$, $\Delta_{u}$ quantifies how separable $a^{\star}(x)$ is from the rest of actions. In other words, the magnitude of $\Delta_{u}$ quantifies the difficulty of the problem. $\Delta_{u}=0$ for all $u\in[0,1)$ means that the optimal action $a^{\star}(x)$ is not distinguishable from the second best action, almost surely. Oppositely, $\Delta_{u}=1$ for a specific $u$ means that with probability 1u1-u, the costs are deterministic and the optimal action $a^{\star}(x)$ is the only one to have nonzero cost (of 1-1). We want to avoid the first case, as the existence of a strictly positive $\Delta_{u}>0$ will be important next. Finally, we introduce the minimum coverage of the optimal actions under $\pi_{k}$:
@@
The proof is given in Appendix˜D. The condition $\Delta_{u}>0$ is analogous to margin assumptions commonly used in multi-arm contextual bandits (Hu et al., 2022; Bibaut et al., 2021), while $C^{\star}_{k}>0$ imposes a coverage requirement ensuring that $\pi_{k}$ sufficiently explores optimal actions. Importantly, this condition is strictly weaker than the uniform coverage assumption imposed in Bibaut et al. (2021); Zenati et al. (2023). Instead of controlling $L$, a similar inequality on the variance of importance weights was used as an assumption to obtain fast rates in Zenati et al. (2023). Here, we prove such result and characterize when it is possible to obtain. For instance, this inequality cannot hold when the costs are constant. Finally, while we formulate the conditions in terms of $\Delta_{u}$ and $C^{\star}_{k}$ for interpretability, a sharper, but less transparent, formulation is given in Sec.˜D.3.
Lemma˜4.3 shows that the expansion of $L$ is controlled by a difference of risks. Smaller $\gamma_{k}$ sharpens this inequality, which is the case for problems with distinguishable optimal actions and near optimal logging policies $\pi_{k}$. Plugging this inequality in Proposition˜4.2 then yields a recursive control on the sequence $\left(R(\pi_{k})\right)_{k\geq 0}$ (see Cor.˜D.5). Such a recursive relation can then be transformed into an explicit accelerated convergence rate for the output of Algorithm˜2, provided the following assumption holds: for all $k\geq 0$, the algorithm outputs policies $\pi_{k}$ such that
$\forall k\geq 0\,,\quad C^{\star}_{k}\geq C^{\star}>0.$
Intuitively, this is plausible since the initial policy $\pi_{0}$ is expected to cover optimal actions, and subsequent policies $\pi_{k}$ for $k\geq 1$ should become progressively closer to $\pi_{\star}$. Formally, this condition is guaranteed by Assumption˜3.2, which requires the policy class to sufficiently cover the optimal actions. Combined with the separability condition ($\Delta_{u}>0$ for some $u\in[0,1)$), we obtain the uniform upper bound on $\gamma_{k}$:
$\forall k,\quad\gamma_{k}\leq\gamma=\frac{3}{\Delta_{u}(1-u)C^{\star}}<+\infty\,.$
With this condition, we state our principal result below.
@@
Let $n_{j}=m>0$ for all $j$. Then, for any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm˜2 at time $k$ with parameter $\lambda_{m}:=\frac{1-\alpha}{8\gamma\sqrt{m}}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
@@
Thm.˜4.4 is proven in Appendix˜D and provides an accelerated rate of $\mathcal{O}\left(\frac{1}{\sqrt{m}k^{\alpha}}\right)$ compared to the $\mathcal{O}\left(\frac{1}{\sqrt{mk}}\right)$ of Algorithm˜1 as long as $\alpha>1/2$. Again, we recover the same rate as Sakhi et al. (2024) in the batch setting (i.e. policy after one batch update: $k=1,m=n$) and improve the rate in the online setting (i.e. updating the behavior policy after each interaction: $k=n,m=1$) to $\mathcal{O}(1/n^{\alpha})$. Thm.˜4.4 also works for any intermediate regime between batch and online settings, and suggests that considering several updates is beneficial to learn efficiently. Indeed, as long as we choose a $\lambda$ that ensures $\alpha>1/2$, the number of updates $k$ will have a stronger influence than the size of the batch $m$. It should be noted, however, that the constant $C_{\alpha}$ deteriorates as $\alpha$ approaches $1$. Larger values of $\lambda_{m}$ yield smaller constants but slower convergence rates, whereas smaller $\lambda_{m}$ yield faster rates at the cost of heavier constants. In practice, this creates a trade-off between accelerating convergence and keeping $C_{\alpha}$ at reasonable levels.
Finally, to capture a broad range of real-world scenarios where the size of the batches received $n_{k}$ may vary over time, we extend these results in Cor.˜4.5, which provides convergence guarantees for Algorithm˜2 in this more general setting.
@@
Let $\beta_{1}:=\sup_{k\geq 0}\frac{k+1}{N_{k}}\leq 1$. Assume there exists $\beta_{2}\in\mathbb{R}^{+}$ such that for all $k$, $n_{k}\leq\beta_{2}$. For any $\alpha\in[0,1)$, denote by $\pi_{k+1}$ the output of Algorithm˜2 at time $k$ with parameter $\lambda=\frac{1}{1+2^{2+\alpha}\gamma\beta_{1}\beta_{2}B(\alpha)}$. Then with probability at least $1-\delta$ over $S$, for all $k\geq 0$:
$R(\pi_{k+1})-R(\pi^{\star})\leq\frac{C_{\alpha}}{(k+1)^{\alpha}},$
where $C_{\alpha}\leq\max\left(1,\frac{\operatorname{KL}(\mathrm{Q}^{\star},P)+\log(2/\delta)}{\lambda^{2}\gamma\beta_{2}B(\alpha)}\right)$, $B(\alpha)=\frac{1}{1-\alpha}$.
Observe that the convergence rate in Cor.˜4.5 is weaker than that of Thm.˜4.4, since it no longer depends explicitly on the batch sizes $n_{j}$. This relaxation is necessary to accommodate adversarial scenarios where some $n_{j}$ may be as small as 1. For example, let $n\geq 2$ and consider a sequence where $n_{2j-1}=m$ and $n_{2j}=1$ for all $j\geq 1$. Then, for all $k\geq 0$, we have $N_{2k-1}=k(m+1)-1$ and $N_{2k}=k(m+1)$. Thus, for all $k$, it holds that $\frac{k+1}{N_{k}}\leq\frac{2}{m}$, and the assumptions of Cor.˜4.5 are satisfied. Importantly, Cor.˜4.5 matches the optimal convergence rate of Thm.˜4.4 in the fully online regime (i.e., $k=n$, $n_{j}=1$). Having established that Algorithm˜2 enjoys provably accelerated convergence rates, we turn to empirical validation to assess whether this theoretical advantage is reflected in practice.
Figure 1: $R(\pi_{k})$ at each intermediate step, for $k=10$.
@@ 5 EXPERIMENTS
General Setup. Our experimental framework follows the standard multiclass-to-bandit conversion widely used in prior work (Dudik et al., 2014; Swaminathan and Joachims, 2015). Each multiclass dataset consists of feature-label pairs, which we transform into contextual bandit problems by treating features as contexts and labels as actions. The reward $r$ for taking action $a$ for context $x$ is modeled as Bernoulli with probability $p_{x}=\epsilon+\mathds{1}\left[a=a^{\star}(x)\right](1-2\epsilon)$, where $a^{\star}(x)$ is the true label of features $x$, and $\epsilon$ is a noise parameter, that we set to $\epsilon=0.2$ in all our experiments. This setup ensures an average reward of $1-\epsilon$ for the optimal action $a^{\star}(x)$ and $\epsilon$ for all others. In our experiments, we use parametrized Linear Gaussian Policies (Sakhi et al., 2023a; Aouali et al., 2023) as our class of policies. These policies interact with the contextual bandit environment to construct a logged bandit feedback dataset in the form $\{x_{i},a_{i},c_{i}\}_{i\in[n]}$, where $c_{i}=-r_{i}$ is the associated cost. Appendix˜E details the experimental design and obtained results.
Effect of Multiple Deployments. Unlike prior batch approaches, our sequential algorithms support multiple updates of the behavior policy $\pi_{0}$, while seamlessly reducing to the batch case when $k=1$. Using the MNIST (LeCun et al., 2010) and CIFAR100 (Krizhevsky, 2009) datasets, we evaluate how the number of intermediate policy updates $k$ affects the performance of the final learned policy $\pi$. Starting from a uniform policy $\pi_{0}$, we run both SeqLS (Alg 1) and the improved SeqAdjLS (Alg 2). For each $k\in\{1,5,10,100\}$, we fix the total number of interactions to $N_{k}=N$, resulting in $m=N/k$ interactions per deployment. The case $k=1$ corresponds to the static setting studied in Sakhi et al. (2024), where all data is collected using $\pi_{0}$. For $k>1$, each intermediate policy $\pi_{j}$ interacts with $m$ examples, collectively producing $N$ sequentially collected samples. We set $\lambda=1/\sqrt{m}$, following our theoretical guidance.
Table˜1 summarizes the results: increasing $k$ consistently improves final policy performance, with diminishing returns as $\pi$ nears optimality (e.g., MNIST at $k=10$ and $100$). The gains are especially notable on CIFAR100, where the larger action space ($|\mathcal{A}|=100$) makes uniform policies particularly suboptimal, thus highlighting the importance of the sequential updates. Finally, our results demonstrate the empirical advantage of SeqAdjLS, validating the need for our refined estimator. Across all settings, and especially the more challenging ones, SeqAdjLS consistently outperforms SeqLS, providing empirical evidence that the adjustment is essential for achieving accelerated convergence.
@@
Dataset $k$ SeqLS (1) SeqAdjLS (2)
MNIST 11 0.674-0.674 0.678-0.678
55 0.758-0.758 0.766-0.766
1010 0.782-0.782 $\mathbf{-0.791}$
100100 0.781-0.781 $\mathbf{-0.789}$
CIFAR100 11 0.197-0.197 0.207-0.207
55 0.344-0.344 0.380-0.380
1010 0.406-0.406 0.455-0.455
100100 0.508-0.508 $\mathbf{-0.558}$
Benchmarks. Having demonstrated the benefits of intermediate updates and the effectiveness of our proposed SeqAdjLS algorithm, we now compare our methods against the recently introduced SCRM approach Zenati et al. (2023), which extends the CRM principle of Swaminathan and Joachims (2015) to the sequential setting. We evaluate all methods on MNIST, FashionMNIST (Xiao et al., 2017), EMNIST (Cohen et al., 2017), and CIFAR100 with varying action space sizes, fixing the total number of interactions to $N$, setting $k=10$, and using theoretically recommended values for $\lambda$ to ensure a fair comparison.
SeqLS and SeqAdjLS divide the interaction budget uniformly across rounds ($n_{j}=N/k$), reusing all previously collected data to update policies. In contrast, SCRM employs exponentially increasing interaction sizes ($n_{j}=n_{0}\cdot 2^{j}$ with $n_{0}=\lceil N/2^{k}\rceil$) and updates policies using only the most recent batch, rather than the cumulative data ($n_{j}$ instead of $N_{j}$). Each method is evaluated across six random seeds, and the results are summarized in Fig.˜1.
We observe that SCRM tends to be overly conservative, particularly in early rounds when the number of interactions is small. Although it improves upon the initial behavior policy $\pi_{0}$ over time, it lags significantly behind both of our sequential methods. While it narrows the gap with SeqLS on CIFAR100, it is consistently outperformed, especially by SeqAdjLS, which achieves superior results across all datasets. These findings further confirm the empirical advantages of our approach. Additional experimental details and results are provided in Appendix˜E.